OpenStudy (anonymous):

What does a dependent system, and how does it work? Could you please show me an example.

OpenStudy (anonymous):

What level of detail do you require? IE for which class?

OpenStudy (anonymous):

for math

OpenStudy (anonymous):

simple explanation such as y=x+2

OpenStudy (anonymous):

an example of a dependent system would be x + y = 1 and 2x+2y=2 Those two equations describes the same line and the solution x,y would be infinite.